# CorMulT: A Semi-supervised Modality Correlation-aware Multimodal Transformer for Sentiment Analysis

### v1

@@
Abstract.
Multimodal sentiment analysis is an active research area that combines multiple data modalities, e.g., text, image and audio, to analyze human emotions and benefits a variety of applications. Existing multimodal sentiment analysis methods can be classified as modality interaction-based methods, modality transformation-based methods and modality similarity-based methods. However, most of these methods highly rely on the strong correlations between modalities, and cannot fully uncover and utilize the correlations between modalities to enhance sentiment analysis. Therefore, these methods usually achieve bad performance for identifying the sentiment of multimodal data with weak correlations. To address this issue, we proposed a two-stage semi-supervised model termed Correlation-aware Multimodal Transformer (CorMulT) which consists pre-training stage and prediction stage. At the pre-training stage, a modality correlation contrastive learning module is designed to efficiently learn modality correlation coefficients between different modalities. At the prediction stage, the learned correlation coefficients are fused with modality representations to make the sentiment prediction. According to the experiments on the popular multimodal dataset CMU-MOSEI, CorMulT obviously surpasses state-of-the-art multimodal sentiment analysis methods.
@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Computing methodologies Machine learning approachesccs: Computing methodologies Philosophical/theoretical foundations of artificial intelligence
1. Introduction
@@
Multimodal sentiment analysis combines multiple perceptual modalities such as text, image, audio and video to analyze and understand human emotions. Currently, multimodal sentiment analysis has been widely applied in various domains, including predicting movie box office performance (Asur and Huberman, 2010), predicting stock market performance (Bol, 2011), and predicting political election results (Tumasjan et al., 2011).
@@
The general paradigm of multimodal sentiment analysis is to extract emotional information from multiple modalities and fuse them together to achieve a more comprehensive and accurate understanding of emotions. Existing methods for multimodal sentiment analysis can be roughly classified into three categories, i.e., modality interaction-based methods, modality transformation-based methods, and modality similarity-based methods. Modality interaction-based methods emphasize on learning the interactions and dependencies between different modalities to combine them together. The representative methods are MultiModal InfoMax (Han et al., 2021) and MMLatch (Paraskevopoulos et al., 2022). Modality transformation-based methods select one modality as the primary modality, and extract information from other modalities to obtain cross-modal representation. The representative methods include BAFN (Tang et al., 2023) and RAVEN (Wang et al., 2019). Modality similarity-based methods compare the features between different modalities for deep understanding of sentiment, and the representative methods include Misa (Hazarika et al., 2020) and MUET (Liu et al., 2022).
@@
However, most existing methods for multimoal sentiment analysis heavily rely on the assumption of strong correlations between modalities, and have not yet fully explored and examined the correlations between modalites. In this case, they may achieve bad performance for identifying the sentiment for multimodal data with weak correlations. In general, strong modality correlations are characterized by a well-aligned, consistent, and strongly related representation of objects across visual, textual, and audio aspects, with minimal noise interference such as irrelevant visual elements and background sounds. Conversely, weak modality correlations are often plagued with different discrepancies as illustrated in Fig. 1. Voice discrepancy arises when non-primary sounds overshadow the main audio subjects. Content discrepancy occurs when there is an inconsistency between the objects described in text and those depicted in images. Alignment discrepancy is the inability to achieve a complete synchronization across modalities. Clarity discrepancy refers to the presence of excessive noises that muddle the primary signals. In practice, it is difficult for existing multimodal sentiment analysis methods to accurately decide the sentiment of data samples with such discrepancies.
@@
Figure 1. The multimodal samples with weak modality correlations, including voice discrepancy, content discrepancy, alignment discrepancy, and clarity discrepancy.
@@
  • We identified the issue of weak modality correlations in existing multimodal sentiment analysis methods, and proposed the CorMulT model to address this issue by accurately learning the correlations between modalities and fusing the learned correlations with modality representation to enhance sentiment analysis.
@@
The remainder of the work is organized as follows. Section 2 reviews the related work on multimodal sentiment analysis. Section 3 presents the proposed CorMulT method in detail. Section 4 reports the experimental results, and Section 5 concludes this work.
2. Related Work
@@
2.1. Modality Interaction-based Methods
@@
These methods focus on capturing the dynamic interactions and dependencies between different modalities during the process of combining information from multiple modalities. Wei et al. (Han et al., 2021) proposed MultiModal InfoMax (MMIM) to maximize the mutual information between unimodal input pairs, and that between the fusion results and low-level unimodal input representations to preserve task-relevant information during multimodal fusion. Paraskevopoulos et al. (Paraskevopoulos et al., 2022) utilized a forward feedback mechanism to capture the top-down cross-modal interactions and achieved multi-modal fusion during network training by masking sensory inputs with higher-level representations extracted from each modality. Kumar et al. (Kumar and Vepa, 2020) utilized self-attention to capture the long-term context and employed a gating mechanism to selectively learn weak contributing features. Sun et al. (Sun et al., 2023) proposed a general debiasing framework based on Inverse Probability Weighting (IPW), which adaptively assigns small weights to the samples with larger bias to avoid fitting the spurious correlations between multimodal features and sentiment labels. Xue et al. (Xue et al., 2023) proposed a multi-level attention map network (MAMN) to filter noise before multimodal fusion and capture the consistent and heterogeneous correlations among multi-granularity features. Quan et al. (Quan et al., 2022) proposed a multimodal comparative learning interaction module to better focus on the characteristics of different modalities. He et al. (He et al., 2023c) proposed a multimodal mutual attention-based sentiment analysis (MMSA) framework that uses three levels of subtasks to preserve the unimodal unique semantics and enhance the common semantics of multimodal data. Wang et al. (Wang et al., 2023a) introduced supervised contrastive learning to learn effective and comprehensive multimodal data representation. Liu et al. (Liu et al., 2023) proposed the Scanning, Attention and Reasoning (SAR) model for multimodal sentiment analysis, where a perceptual scanning model is designed to perceive the image and text content, as well as the intrinsic correlations between them. Wang et al. (Wang et al., 2023b) proposed a model to assist in learning modality representations with multitask learning and contrastive learning. Tsai et al. (Tsai et al., 2019) repeatedly reinforce the features of the target modality using low-level features of the source modality, thus ensuring cross-modal interactions and addressing the alignment issue between different modalities. Yu et al. (Yu et al., 2023) proposed a cyclic memory enhancement network for capturing the long-term dependencies in multimodal data. Ping et al. (He et al., 2023a) proposed a novel temporal position prediction task for speech-text alignment to deal with the cross-modal alignment issue. Cheng et al. (Cheng et al., 2023) and Ye et al. (He et al., 2022) proposed Attentional Temporal Convolutional Network (ATCN) and Multimodal Temporal Attention (MMTA) to extract unimodal temporal features for adaptive inter-modal balance. Lin et al. (Lin and Hu, 2023) and He et al. (He et al., 2023b) narrowed the gap among different modalities by transferring the sentiment-related knowledge.
2.2. Modality Transformation-based Methods
@@
In this category of methods, one modality serves as the primary modality while the others serve as supplementary modalities. For example, Khan et al. (Khan and Fu, 2021) developed a two-stream model that converts image modality to textual representations as auxiliary text, thereby providing more information for the text modality. Wang et al. (Wang et al., 2019) obtained the text drift vectors using non-textual data, e.g., audio and image modalities, thus achieving drifted word embedding. Majumder et al. (Majumder et al., 2017) mapped the personality traits from textual documents to audio and visual modalities to improve the effectiveness of sentiment analysis. Tang et al. (Tang et al., 2023) introduced a more discriminating text modality to guide the dynamic learning of emotional context within modalities, thus reducing the redundant information in auxiliary modalities during the modality transformation process. He et al. (He and Hu, 2022) fused nonverbal embedding with language embedding before inputting them into Bert to obtain multimodal representations. However, according to Gan et al. (Gan et al., 2015), pure modality transformation methods often underperform the methods that consider the interaction between modalities, which could be because modality transformation methods overlook the inherent correlations between modalities and result in the loss of information.
2.3. Modality Similarity-based Methods
@@
These methods compare the features across different modalities to facilitate the analysis of their sentiment. For instance, Yu et al. (Yu et al., 2021) jointly trained multimodal and unimodal tasks to learn the consistency and differences in modality representations. Hazarika et al. (Hazarika et al., 2020) projected each modality into two subspaces, i.e., modality-invariant subspace and modality-specific subspace, where modality-invariant subspace facilitates cross-modal representation learning to reduce modality discrepancies, and modality-specific subspace captures the distinctive features of each modality. Mai et al. (Mai et al., 2022) proposed curriculum learning for weakly supervised modality correlations, which leverage limited or noisy labeled data to align and correlate features across different modalities, to learn a more discriminative embedding space for multimodal data. However, they often struggle with the ambiguity and sparsity of weak labels, which can lead to suboptimal learning of true modality correlations. In summary, although these methods are capable of learning certain correlations between modalities, they often fall short in further refining and integrating the learned correlations for sentiment analysis, and are thus unable to achieve outstanding performance.
2.4. Discussion
@@
As described above, most existing multimodal sentiment analysis models rely on the assumption of strong correlations between modalities. However, a large amount of multimodal data may not exhibit strong correlations. As a result, the correlations between modalities have not been fully explored and utilized in most models. Besides, it is labor-intensive and impractical to manually filter all the data samples with low modality correlations, and existing methods like cross-modal alignment also cannot solve the problem. Therefore, this work proposed the two-stage model CorMulT. At the pre-training stage, CorMulT learns the modality correlation coefficients via a modality correlation contrastive learning module. At the prediction stage, the learned modality correlation coefficients are fused with the modality representations to enhance the sentiment prediction.
3. Methodology
3.1. Overview
@@
Figure 2. The framework of CorMulT Model which consists of modality correlation contrastive learning module and correlation-aware multimodal transformer module.
3.2. Modality Correlation Contrastive Learning
Modality correlation contrastive learning module utilizes the contrastive learning framework to learn the correlation coefficients between different modalities. This module consists of two submodules, i.e., modality-specific feature extraction and modality correlation evaluation. The modality-specific feature extraction module is responsible for extracting unimodal features from the three modalities, i.e., audio, image and text. The modality correlation evaluation module analyzes the correlations between different modalities.
3.2.1. Modality-Specific Feature Extraction
@@
For audio encoding, we employ the Mel spectrogram (Yang et al., 2023), a widely used visual representation in audio processing that transforms the audio signals into an image-like format. Initially, we capture the spectral characteristics of the audio signals using the Short-Time Fourier Transform (STFT) which decomposes the audio signals into frequency components over short time intervals and produces a time-frequency representation, i.e.,
(1) $X_{\text{stft}}(n,k)=\sum_{m=-\infty}^{\infty}x(m)w(m-n)e^{-j2\pi km/N}$
@@
(2) $H_{m}(k)=\begin{cases}0&\text{if }f_{k}<f_{m-1}\\ \frac{f_{k}-f_{m-1}}{f_{m}-f_{m-1}}&\text{if }f_{m-1}\leq f_{k}<f_{m}\\ \frac{f_{m+1}-f_{k}}{f_{m+1}-f_{m}}&\text{if }f_{m}\leq f_{k}<f_{m+1}\\ 0&\text{otherwise}\end{cases}$
@@
(3) $X_{\text{mel}}(n,m)=\log\left(\sum_{k=0}^{N-1}|X_{\text{stft}}(n,k)|^{2}H_{m}(k)\right)$
where $X_{\text{mel}}(n,m)$ denotes the log-transformed Mel spectrogram value for the $m^{th}$ Mel frequency band at the $n^{th}$ time frame. The Mel spectrogram matrix is defined in the context of $F_{A}\subseteq\mathbb{R}^{b\times t\times m}$, where $b$ represents the batch size, $t$ represents the number of time frames, and $m$ represents the number of Mel frequency bins. This logarithmic scaling of the Mel spectrogram not only enhances the representation of the audio signals by emphasizing the perceptually important aspects of the sound, but also stabilizes the numerical range, thus facilitating more effective model training.
@@
For text encoding, the input text $T$ is first tokenized using the BertTokenizer to generate a sequence of tokens $W$. Next, a vocabulary $V$ is created using the CMU_MOSEI dataset, which contains a collection of words and their corresponding numerical representations. The tokens $W$ are then encoded using the vocabulary $V$ to obtain the encoded sequence $S$. To ensure uniformity in the input size, the encoded sequence $S$ is padded or truncated to a pre-defined sequence length. The final representation of the text input is $F_{T}\subseteq\mathbb{R}^{b\times s}$, where $b$ is the batch size and $s$ is the sequence length.
@@
3.2.2. Sample Generation for Contrastive Learning

### v2

# CorMulT: A Modality Correlation-aware Multimodal Transformer for Sentiment Analysis

@@
Abstract
Multimodal sentiment analysis is an active research area that combines multiple data modalities, e.g., text, image and audio, to analyze human emotions, and benefits a variety of applications. Existing multimodal sentiment analysis methods can be roughly classified as modality interaction-based methods, modality transformation-based methods and modality similarity-based methods. However, most of these methods highly rely on the strong correlations between modalities, and cannot fully uncover and utilize the correlations between modalities to enhance sentiment analysis. Therefore, these methods usually achieve unsatisfactory performance for identifying the sentiment of multimodal data with weak correlations. To address this issue, we proposed a two-stage semi-supervised model termed Correlation-aware Multimodal Transformer (CorMulT) which consists of pre-training stage and prediction stage. At the pre-training stage, a modality correlation contrastive learning module is designed to efficiently learn modality correlation coefficients between different modalities. At the prediction stage, the learned correlation coefficients are fused with modality representations to make the sentiment prediction. According to the experiments on the popular multimodal dataset CMU-MOSEI, CorMulT obviously surpasses the state-of-the-art multimodal sentiment analysis methods.
Index Terms:
@@ Abstract
I Introduction
@@
Multimodal sentiment analysis combines multiple perceptual modalities such as text, image, and audio to analyze human emotions. Currently, multimodal sentiment analysis has been widely applied in various domains, including predicting movie box office performance [1], predicting stock market performance [2], and predicting political election results [3].
@@
The general paradigm of multimodal sentiment analysis is to extract emotional information from multiple modalities and fuse them together to achieve a comprehensive and accurate understanding of emotions. Existing methods for multimodal sentiment analysis can be roughly classified into three categories, i.e., modality interaction-based methods, modality transformation-based methods, and modality similarity-based methods. Modality interaction-based methods emphasize on learning the interactions and dependencies between different modalities to combine them together. The representative methods are MultiModal InfoMax [4] and MMLatch [5]. Modality transformation-based methods select one modality as the primary modality, and extract information from other modalities to obtain cross-modal representation. The representative methods include BAFN [6] and RAVEN [7]. Modality similarity-based methods compare the features of different modalities for deep understanding of sentiment, and the representative methods include Misa [8] and MUET [9].
@@
However, most existing methods for multimoal sentiment analysis heavily rely on the assumption of strong correlations between modalities, and have not yet fully explored and examined the correlations between modalites. In this case, they may achieve unsatisfactory performance for identifying the sentiment for multimodal data with weak correlations. In general, strong modality correlations are characterized by a well-aligned, consistent, and strongly related representation of objects across visual, textual, and audio aspects, with minimal noise interference such as irrelevant visual elements and background sounds. Conversely, weak modality correlations are often plagued with different discrepancies as illustrated in Fig. 1. Voice discrepancy arises when non-primary sounds overshadow the main audio subjects. Content discrepancy occurs when there is an inconsistency between the objects described in text and those depicted in images. Alignment discrepancy is the inability to achieve a complete synchronization across modalities. Clarity discrepancy refers to the presence of excessive noises that muddle the primary signals. In practice, it is difficult for existing multimodal sentiment analysis methods to accurately decide the sentiment of data samples with such discrepancies.
@@
Figure 1: The multimodal samples with weak modality correlations, including voice discrepancy, content discrepancy, alignment discrepancy, and clarity discrepancy.
@@
  • We identified the issue of weak modality correlations in multimodal sentiment analysis, and proposed the CorMulT model to address this issue by accurately learning the correlations between modalities and fusing the learned correlations with modality representation to enhance sentiment analysis.
@@
The remainder of the work is organized as follows. Section II reviews the related work on multimodal sentiment analysis. Section III presents the proposed CorMulT method in detail. Section IV reports the experimental results, and Section V concludes this work.
II Related Work
@@
II-A Modality Interaction-based Methods
These methods focus on capturing the dynamic interactions and dependencies between different modalities when combining information from multiple modalities. Wei et al. [4] proposed MultiModal InfoMax (MMIM) to maximize the mutual information between unimodal pairs, and that between the fusion results and low-level unimodal representations to preserve task-relevant information during multimodal fusion. Paraskevopoulos et al. [5] utilized a forward feedback mechanism to capture the top-down cross-modal interactions and achieved multimodal fusion by masking sensory inputs with higher-level representations extracted from each modality. Kumar et al. [10] utilized self-attention to capture the long-term context and employed a gating mechanism to selectively learn weak contributing features.
@@ II-A Modality Interaction-based Methods
Sun et al. [11] proposed a general debiasing framework based on Inverse Probability Weighting (IPW), which adaptively assigns small weights to the samples with larger bias to avoid fitting the spurious correlations between multimodal features and sentiment labels. Xue et al. [12] proposed a multi-level attention map network (MAMN) to filter noise before multimodal fusion and capture the consistent and heterogeneous correlations among multi-granularity features. Quan et al. [13] proposed a multimodal comparative learning interaction module to better focus on the characteristics of different modalities. He et al. [14] proposed a multimodal mutual attention-based sentiment analysis (MMSA) framework that uses three levels of subtasks to preserve the unimodal unique semantics and enhance the common semantics of multimodal data. Wang et al. [15] introduced supervised contrastive learning to learn effective and comprehensive multimodal data representation. Liu et al. [16] proposed the Scanning, Attention and Reasoning (SAR) model for multimodal sentiment analysis, where a perceptual scanning model is designed to perceive the image and text content, and the intrinsic correlations between them. Wang et al. [17] proposed a model to assist in learning modality representations with multitask learning and contrastive learning.
Tsai et al. [18] repeatedly reinforce the features of the target modality using low-level features of the source modality, thus ensuring cross-modal interactions and addressing the alignment issue between different modalities. Yu et al. [19] proposed a cyclic memory enhancement network for capturing the long-term dependencies in multimodal data. Ping et al. [20] proposed a novel temporal position prediction task for speech-text alignment to deal with the cross-modal alignment issue. Cheng et al. [21] and Ye et al. [22] proposed Attentional Temporal Convolutional Network (ATCN) and Multimodal Temporal Attention (MMTA) to extract unimodal temporal features for adaptive inter-modal balance. Lin et al. [23] and He et al. [24] narrowed the gap among different modalities by transferring the sentiment-related knowledge.
II-B Modality Transformation-based Methods
@@
In this category of methods, one modality serves as the primary modality while the others serve as supplementary modalities. For example, Khan et al. [25] developed a two-stream model that converts image modality to textual representations as auxiliary text, thereby providing more information for the text modality. Wang et al. [7] obtained the text drift vectors using non-textual data, e.g., audio and image modalities, thus achieving drifted word embedding. Majumder et al. [26] mapped the personality traits from textual documents to audio and visual modalities to improve the effectiveness of sentiment analysis. Tang et al. [6] introduced a more discriminating text modality to guide the dynamic learning of emotional context within modalities, thus reducing the redundant information in auxiliary modalities during the modality transformation process. He et al. [27] fused non-verbal embedding with language embedding before inputting them into Bert to obtain multimodal representations. However, according to Gan et al. [28], pure modality transformation methods often underperform the methods that consider the interaction between modalities because modality transformation methods may overlook the inherent correlations between modalities and result in information loss.
II-C Modality Similarity-based Methods
@@
These methods compare the features across different modalities to facilitate the analysis of their sentiment. For instance, Yu et al. [29] jointly trained multimodal and unimodal tasks to learn the consistency and difference in modality representations. Hazarika et al. [8] projected each modality into two subspaces, i.e., modality-invariant subspace and modality-specific subspace, where modality-invariant subspace facilitates cross-modal representation learning to reduce modality discrepancies, and modality-specific subspace captures the distinctive features of each modality. Mai et al. [30] proposed curriculum learning for weakly supervised modality correlations, and leveraged limited or noisy labeled data to align features across different modalities, thus learning a more discriminative embedding space for multimodal data. However, they often struggle with the ambiguity and sparsity of weak labels, which may lead to suboptimal learning of true modality correlations. In summary, although these methods are capable of learning certain correlations between modalities, they often fall short in further integrating the learned correlations for sentiment analysis, and are thus unable to enhance the analysis performance.
II-D Discussion
@@
As described above, most existing multimodal sentiment analysis models rely on the assumption of strong correlations between modalities. However, as discussed in Section I, a large amount of multimodal data may not exhibit strong correlations. As a result, the correlations between modalities have not been fully explored and utilized in most models. Besides, it is labor-intensive and impractical to manually filter all the data samples with low modality correlations, and existing methods like cross-modal alignment also cannot solve the problem. Therefore, this work proposed the two-stage model CorMulT. At the pre-training stage, CorMulT learns the modality correlation coefficients via a modality correlation contrastive learning module. At the prediction stage, the learned modality correlation coefficients are fused with the modality representations to enhance the sentiment prediction.
III Methodology
III-A Overview
@@
Figure 2: The framework of CorMulT Model which consists of modality correlation contrastive learning module and correlation-aware multimodal transformer.
Note: The CorMulT model initiates with the Modality Correlation Evaluation (MCE) module using contrastive learning to determine modality correlation coefficients (Stage 1). The trained MCE is integral for enabling the Multimodal Transformer to understand and utilize the relationships between different modalities for more accurate predictions (Stage 2).
III-B Modality Correlation Contrastive Learning
Modality correlation contrastive learning module utilizes the contrastive learning framework to learn the correlation coefficients between different modalities. This module consists of two submodules, i.e., modality-specific feature extraction and modality correlation evaluation. The modality-specific feature extraction is responsible for extracting unimodal features from the three modalities, i.e., audio, image and text. The modality correlation evaluation module analyzes the correlations between different modalities.
III-B1 Modality-Specific Feature Extraction
@@
For audio encoding, we employ the Mel spectrogram [31] which is a widely used visual representation in audio processing that transforms the audio signals into an image-like format. Initially, we capture the spectral characteristics of the audio signals using the Short-Time Fourier Transform (STFT) which decomposes the audio signals into frequency components over short time intervals and produces a time-frequency representation, i.e.,
$X_{\text{stft}}(n,k)=\sum_{m=-\infty}^{\infty}x(m)w(m-n)e^{-j2\pi km/N}$ (1)
@@
$H_{m}(k)=\begin{cases}0&\text{if }f_{k}<f_{m-1}\\ \frac{f_{k}-f_{m-1}}{f_{m}-f_{m-1}}&\text{if }f_{m-1}\leq f_{k}<f_{m}\\ \frac{f_{m+1}-f_{k}}{f_{m+1}-f_{m}}&\text{if }f_{m}\leq f_{k}<f_{m+1}\\ 0&\text{otherwise}\end{cases}$ (2)
@@
$X_{\text{mel}}(n,m)=\log\left(\sum_{k=0}^{N-1}|X_{\text{stft}}(n,k)|^{2}H_{m}(k)\right)$ (3)
where $X_{\text{mel}}(n,m)$ denotes the log-transformed Mel spectrogram value for the $m^{th}$ Mel frequency band at the $n^{th}$ time frame. The Mel spectrogram matrix $X_{\text{mel}}$ is an element of the set $F_{A}$, which is a subset of $\mathbb{R}^{b\times t\times m}$, where $b$ represents the batch size, $t$ denotes the number of time frames, and $m$ represents the number of Mel frequency bands. This logarithmic scaling of the Mel spectrogram not only enhances the representation of the audio signals by emphasizing the perceptually important aspects of the sound, but also stabilizes the numerical range, thus facilitating more effective model training.
@@
For text encoding, the input text $T$ is first tokenized using the BertTokenizer to generate a sequence of tokens $W$. Next, a vocabulary $V$ is created using the CMU_MOSEI dataset, and $V$ contains a collection of words and their corresponding numerical representations. The tokens $W$ are then encoded using the vocabulary $V$ to obtain the encoded sequence $S$. To ensure uniformity in the input size, the encoded sequence $S$ is padded or truncated to a pre-defined sequence length. The final representation of the text input is $F_{T}\subseteq\mathbb{R}^{b\times s}$, where $b$ is the batch size and $s$ is the sequence length.
@@
III-B2 Sample Generation for Contrastive Learning